Suppose a heat engine has an efficiency of 0.177 and produces 5.40×103 J of work.

How much energy is pulled in from the hot reservoir?

How much energy is rejected by the engine? (This energy is "exhausted" to the cold reservoir.))
Answer #1

energy is pulled in from the hot reservoir IS,

Q = W/e = 5.40×103 J/ 0.177 = 30508.47 J

energy rejected by the engine is,

E = 30508.47 J - 5.40×103 J = 25108.47 J